# Evaluating Decision Optimality of Autonomous Driving via Metamorphic Testing

### v1

@@ 4.2. MR Check $\epsilon$
Figure 4. MR checking with grid map. The above are the driving paths and the below are the locations of the corresponding paths in the grid map.
@@
Figure 4 illustrates the basic idea of our method. We define an optimal area around the optimal path $\tau(s)$ in terms of the grid map. Any driving path falling or mostly falling within this area is considered optimal and satisfies the MR. For example, the green area in Figure 4(a) defines the optimal area of the driving path in the ODS. The driving path in Figure 4(b) lies in the optimal area, resulting in the highest similarity, and thus meets the MR criterion, while most of the path in Figure 4(c) (i.e., the red area) falls outside the optimal area, resulting in a low similarity, and the MR is violated.
@@
Specifically, the comparison between the driving path $\tau(s)$ and the driving path $\tau(s^{\prime})$ is implemented by a grid-based approach, as shown in Figure 4. The map is divided into grids, and each driving path is mapped to a set of grids. Technically, a driving path of the ego vehicle is represented as a sequence of locations $\tau(s)=\{p_{0}^{0},\ldots,p_{k}^{0}\}$, where $p_{i}^{0}$ is the position at frame $i$. To collect the covered grids, a function $g$ is used to map each position $p$ to the grid where it is located. The covered grids for the path $\tau(s)$ can be denoted as:
(5) $C_{\tau(s)}=\{{g}(p)\,|\,p\in\tau(s)\}$
@@
To check the MR, the similarity between the covered grids of $\tau(s)$ and $\tau(s^{\prime})$ is computed. The MR checking is defined using a predefined threshold $\varepsilon$ as follows:
(6) $\epsilon(\tau(s),\tau(s^{\prime}))=\frac{|C_{\tau(s)}\cap C_{\tau(s^{\prime})}|}{|C_{\tau(s)}\cup C_{\tau(s^{\prime})}|}>\varepsilon$
@@
If the similarity between the covered grids is greater than the threshold $\varepsilon$, the MR is satisfied and $s^{\prime}$ is recognized as an ODS.
4.3. Feedback $\mathcal{D}$
@@
To guide the search of NoDSs, a feedback is necessary to select high-quality individuals from the candidate population (Line 1 of Algorithm 1). The grid similarity $\epsilon(\tau(s)$, $\tau(s^{\prime}))$ (in Equation 6) provides a direct choice for the feedback. However, the calculation of grid similarity is relatively coarse-grained as it only considers the spatial perspective of the ego vehicle’s motion. While this coarse-grained calculation is beneficial for MR checking, as it considers different possible optimal paths, it may not be as specific and effective for guiding the testing process to generate NoDSs (see the evaluation results in Section 5.2.2). To address this, another fitness is proposed that incorporates more fine-grained feedback from both the driving path and the behavior of the ego vehicle. The former focuses on the spatial characteristics of the ego vehicle’s motion, while the latter more on the temporal ones.
@@
Driving Path Feedback. The main objective of Decictor is to maximize the difference between the driving paths such that MR is not satisfied. Considering the potential different lengths of $\tau(s)$ and $\tau(s^{\prime})$, we use the shortest distance between $\tau(s)$ and $\tau(s^{\prime})$ in the spatial space to measure their difference. A larger distance indicates a higher likelihood that the MR will be violated. The distance is calculated based on the point-wise distance:
(7) $f_{p}(s,s^{\prime})=\frac{1}{n_{s^{\prime}}}\sum_{i=1}^{n_{s^{\prime}}}\min(\{\|p_{i}-p\|\,|\,\forall p\in\tau(s)\})$
@@
Behavior Feedback. In some cases, optimizing only the driving path feedback might be insufficient (see the experimental results in Section 5.2.2), so an additional feedback mechanism is provided, which relies on the behavior of the ego vehicle. This behavior feedback considers factors such as the ego’s velocity, acceleration, and heading. By analyzing the ego’s behavior, it becomes possible to gain insights into how slight changes in behavior could lead to variations in the driving path and increase the overall difference between $\tau(s)$ and $\tau(s^{\prime})$. For example, if the heading of the ego vehicle changes slightly, it may not directly cause a significant increase in driving path differences. However, such a change could serve as a valuable indicator, as altering the heading or acceleration of the ego vehicle could lead to adjustments in the driving path, potentially resulting in an increased difference between paths.
To implement this behavior feedback, the ego’s behavior is collected from its waypoints in the scenario $s$, denoted as
$\mathbf{X}_{s}=((v_{0},a_{0},\theta_{0}),(v_{1},a_{1},\theta_{1}),\ldots,(v_{k},a_{k},\theta_{k}))$
where $v_{i}$, $a_{i}$, and $\theta_{i}$ represent the velocity, acceleration, and heading at timestamp $i$, respectively. To compare the behavior differences between two scenarios $s$ and $s^{\prime}$, the Maximum Mean Discrepancy (MMD) is used as the measure of distance between their behavior distributions. The MMD is a widely used statistical metric for comparing distributions and can effectively quantify differences between two sets of data. The behavior feedback function $f_{b}(s,s^{\prime})$ is defined as follows:
(8) $f_{b}(s,s^{\prime})=f_{MMD}(\mathbf{X}_{s},\mathbf{X}_{s^{\prime}}).$
Finally, the fitness of $s^{\prime}$ in relation to the ODS $s$ is calculated as:
(9) $\mathcal{D}(\tau(s),\tau(s^{\prime}))=f_{p}(s,s^{\prime})+f_{b}(s,s^{\prime}).$
5. EMPIRICAL EVALUATION
Table 1. Comparison results with baselines.
Method %Mutationnormal-↑\,$\uparrow$ #NoDS (#NoDS-Hum)\,$\uparrow$
S1 S2 S3 S4 Avg. S1 S2 S3 S4 Sum
AVFuzzer 41.5 40.5 70.6 10.6 40.8 0.6 (0.3) 1.5 (1.0) 0.0 (0.0) 1.2 (0.5) 3.3 (1.8)
SAMOTA 58.6 50.1 49.6 55.7 53.4 1.5 (0.9) 0.4 (0.3) 0.0 (0.0) 3.3 (1.2) 5.2 (2.4)
BehAVExplor 29.7 36.3 36.6 32.1 33.6 2.5 (1.7) 1.6 (0.7) 0.0 (0.0) 6.3 (2.3) 10.4 (4.7)
Random 73.2 73.1 70.2 76.7 73.3 0.3 (0.0) 3.0 (1.5) 0.2 (0.2) 1.9 (0.7) 5.4 (2.4)
Random-$\delta$ 96.4 97.4 99.0 96.0 97.2 3.0 (1.4) 5.1 (4.6) 4.4 (4.4) 7.2 (2.0) 19.7 (12.4)
Decictor 98.0 97.5 99.4 97.5 98.1 9.4 (3.9) 11.9 (8.6) 9.0 (9.0) 15.7 (6.7) 46.0 (28.2)
@@ 5. EMPIRICAL EVALUATION
Environment. We implement Decictor with Baidu Apollo 7.0 (Baidu, 2019) and its built-in simulation environment SimControl. Baidu Apollo 7.0 is an open-source and industrial-level ADS that supports a wide variety of driving supports.
@@
Driving Scenarios. We evaluate Decictor on a real-world Map Sunnyvale Loop, provided as part of Baidu Apollo. Similar to (Li et al., 2020; Cheng et al., 2023), we manually choose four basic representative scenarios and build the initial ODSs, the inputs of Algorithm 1. These basic scenarios have been widely tested in previous literature (Li et al., 2020; Cheng et al., 2023; Thorn et al., 2018; Zhou et al., 2023; Haq et al., 2022). As shown in the left column of Figure 6, the selected ODSs include left turn ( $S1$), right turn (S2), lane following (S3) and U-turn (S4). All initial ODSs are manually confirmed.
@@
Baselines. In this paper, we selected five baseline methods including Random, Random-$\delta$, AVFuzzer (Li et al., 2020), SAMOTA (Haq et al., 2022), and BehAVExplor (Cheng et al., 2023). Specifically, we implemented two Random strategies: 1) Random, which does not have any feedback and mutation constraints. It randomly generates scenarios by adding or removing dynamic vehicles or static obstacles; 2) Random-$\delta$, which does not have feedback but uses our non-invasive mutation $\delta$. To the best of our knowledge, this is the first paper to specifically evaluate the optimality of decision-making, and there are no direct baselines. Consequently, we understand that comparisons with the three other baselines, primarily designed for identifying safety-critical violations, may not be entirely fair. However, while not their primary focus, they can still generate scenarios with suboptimal decisions. Comparison with these tools can still underscore the need for specialized tools such as Decictor to test the optimality of decisions. Note that these three baselines are originally evaluated on the simulator LGSVL (Rong et al., 2020). Due to the official sunset of LGSVL (LG Electronics, [n. d.]), we customize them to our simulation environment (i.e., SimControl + Apollo) for comparison.
Metrics. To facilitate comparison with the baseline techniques, we have incorporated our MR into them to gather the generated NoDSs. In our experiments, we utilize the metrics #NoDS and #NoDS-Hum to assess the effectiveness of NoDS generation. #NoDS quantifies the number of potential NoDSs, identified through our MR checking with a predetermined threshold $\varepsilon$. To address the inherent subjectivity in determining (non-)optimal decisions, we introduce #NoDS-Hum, which represents the number of NoDSs validated by human assessment. Here, the co-authors independently verify whether each NoDS is truly a non-optimal decision. A NoDS is only counted in #NoDS-Hum if all authors unanimously recognize it as non-optimal, indicating a clear-cut case of a non-optimal decision. It’s important to note that scenarios excluded from #NoDS-Hum don’t necessarily indicate they are still optimal decisions; rather, it means at least one participant was uncertain about the non-optimality. The scenarios not adhering to the non-invasive mutation will be filtered directly, owing to the absence of verifiable ground truth. Moreover, we also assess the success rate of non-invasive mutations, denoted as %Mutation. The evaluation involves replaying the path of the ego vehicle in each mutated scenario alongside the optimal path of the ODS and verifying the completion of the motion task.
@@
Implementation. Following AVFuzzer (Li et al., 2020), we set the population size $N$ in Decictor to $4$. In our mutation, we consider both dynamic vehicles and static obstacles (i.e., traffic cones). The time step $\Delta t$ in the non-invasive mutation is set to 2s. In MR check, the equality threshold $\varepsilon$ is empirically set to $0.6$, and the grid size is set to 2 meters. Similar to previous works (Haq et al., 2022, 2023), we repeat each experiment 10 times and report statistics over these runs. For each run, we used the same budget of four hours, as we found that it was long enough to compare based on our preliminary evaluation.
Figure 5. Comparison of #NoDS for different tools.
5.1. RQ1: Effectiveness of Decictor
Figure 6. Samples of NoDSs detected by Decictor. The red vehicle is the AV. Optimal paths are indicated in green, while non-optimal paths are displayed in red. The first column is the initial ODSs, the second column shows the corresponding NoDSs, and the last column depicts the reproduced scenarios validating that the original optimal paths can be traversed in the NoDSs.
5.1.1. Comparative Results.
Table 1 compares %Mutation, #NoDS, and #NoDS-Hum across four initial ODSs: S1, S2, S3, and S4, and Figure 5 shows the distribution of #NoDS, where the median and the average are represented by an orange bar and a green triangle, respectively. From the results, we can find that Decictor outperforms the baselines in #NoDS and #NoDS-Hum. In detail, on average of #NoDS, Decictor outperforms the best baseline (i.e., Random-$\delta$): S1 (9.4 vs. 3.0), S2 (11.9 vs. 5.1), S3 (9.0 vs. 4.4), and S4 (15.7 vs. 7.2). Among all the detected NoDSs by Decictor, we finally identified 3.9, 8.6, 9.0, and 6.7 NoDS-Hum in S1, S2, S3, and S4, respectively. Even with very rigorous manual filtering, Decictor still significantly outperforms the best baseline (Random-$\delta$). The results indicate the effectiveness of Decictor in identifying NoDSs and NoDS-Hum. We further observed that Random-$\delta$ outperforms both the Random (19.7 vs. 5.4 for the total number of NoDSs) and other three safety-oriented baselines (19.7 vs. 3.3, 5.2, and 10.4 in Sum of #NoDS). This is attributed to the effectiveness of the non-invasive mutation used in Random-$\delta$ and Decictor. Regarding %Mutation (Avg. column), Table 1 reveals that Random and safety-oriented baselines produce a maximum of 73.3% valid mutations (i.e., the original optimal path is not affected), which is significantly lower than non-invasive mutation methods (97.2% for Random-$\delta$ and 98.1% for Decictor on average). Therefore, they are not effective in detecting NoDSs. The comparison of results between Random-$\delta$ and Decictor highlights the effectiveness of the feedback mechanism in Decictor. BehAVExplor performs better than the other two safety-guided baselines due to its diversity feedback mechanism, enabling it to generate scenarios with a wide range of behaviors. As a result, it has a higher likelihood of generating NoDSs.

### v3

@@ IV-B Consistency Check $\epsilon$
Figure 4: Consistency Checking. Above: Driving paths. Below: Locations of the corresponding paths in the grid map.
Given the seed ODS $s^{*}$ and a mutated scenario $s^{\prime}$, we require a criterion $\epsilon$ to determine whether $s^{\prime}$ remains the optimal path in $s^{*}$. A direct approach might be checking if the driving path in $s^{\prime}$ (i.e., $\tau(s^{\prime})$) and the original path of $s^{*}$ (i.e., $\tau(s^{*})$) are the same. However, even in real-world scenarios, human drivers do not strictly adhere to a straight line but exhibit some degree of zig-zag motion, as shown in Fig. 4(b). This motion can still be considered optimal even if it slightly deviates from the original path $\tau(s^{*})$. Thus, we propose an abstraction-based method to quantify the similarity of the two driving paths.
@@
Fig. 4 illustrates the basic idea of our method. We define an optimal area around the optimal path $\tau(s^{*})$ in terms of the grid map. Any driving path falling or mostly falling within this area is considered optimal. For example, the orange area in Fig. 4(a) defines the optimal area of the driving path in the seed ODS. The driving path in Fig. 4(b) does not strictly adhere to the optimal path but remains within the optimal area, achieving the highest similarity and thus meeting the optimality criterion. In contrast, most of the path in Fig. 4(c) (i.e., the blue area) falls outside the optimal area, resulting in a low similarity and a non-optimal path.
@@
Specifically, the comparison between the driving paths $\tau(s^{*})$ and $\tau(s^{\prime})$ is implemented by a grid-based approach, as shown in Fig. 4. The map is divided into grids, and each driving path is mapped to a set of grids. Technically, a driving path of the ego vehicle in a scenario $s$ is represented as a sequence of locations $\tau(s)=\{p_{0}^{0},\ldots,p_{k}^{0}\}$, where $p_{i}^{0}$ is the position at frame $i$. To collect the covered grids, a function $g$ is used to map each position $p$ to the grid where it is located. The covered grids for the path $\tau(s)$ can be denoted as:
$C_{\tau(s)}=\{{g}(p)\,|\,p\in\tau(s)\}$ (5)
@@
To check the decision consistency, the similarity between the covered grids of $\tau(s^{*})$ and $\tau(s^{\prime})$ is computed. The consistency checking is defined using a predefined threshold $\varepsilon$ as follows:
$\epsilon(\tau(s^{*}),\tau(s^{\prime}))=\frac{|C_{\tau(s^{*})}\cap C_{\tau(s^{\prime})}|}{|C_{\tau(s^{*})}\cup C_{\tau(s^{\prime})}|}>\varepsilon$ (6)
@@
If the similarity between the covered grids is greater than the threshold $\varepsilon$, $s^{\prime}$ is recognized as an ODS.
IV-C Feedback $\mathcal{D}$
@@
To guide the search of NoDSs, a feedback is necessary to select high-quality individuals from the candidate population (Line 1 of Algorithm 1). The grid similarity $\epsilon(\tau(s^{*})$, $\tau(s^{\prime}))$ (in Equation 6) provides a direct choice. However, the calculation of grid similarity is relatively coarse-grained as it only considers the spatial perspective of the ego vehicle’s motion. While this coarse-grained calculation is beneficial for consistency checking, it may not be as specific and effective for guiding the testing to generate NoDSs (see our evaluation results in Section V-B2). To mitigate this, another metric is proposed that incorporates more fine-grained feedback calculated from scenario observations (detailed in Section II-B), taking into account the ego vehicle’s driving path and motion behavior. The former focuses on the spatial characteristics of the ego vehicle’s motion, while the latter more on the temporal ones.
@@
Driving Path Feedback. The main objective of Decictor is to maximize the difference between the driving paths such that the non-optimal decision is identified. Considering the potential different lengths of $\tau(s^{*})$ and $\tau(s^{\prime})$, we use the average point-wise distance between $\tau(s^{*})$ and $\tau(s^{\prime})$ in the spatial space to measure their difference, which is calculated as:
$f_{p}(s^{*},s^{\prime})=\frac{1}{n_{s^{\prime}}}\sum_{i=1}^{n_{s^{\prime}}}\min\{\|p_{i}-p\|\,|\,\forall p\in\tau(s^{*})\}$ (7)
@@
Behavior Feedback. In some cases, optimizing only the driving path feedback is difficult, thus an additional feedback mechanism is provided, which is based on the fine-grained behavior of the ego vehicle. This behavior feedback considers factors such as the ego’s velocity, acceleration, and heading. By analyzing the ego’s behavior, it becomes possible to gain insights into how slight changes in behavior could lead to variations in the driving path and increase the overall difference between $\tau(s^{*})$ and $\tau(s^{\prime})$. For example, if the heading of the ego vehicle changes slightly, it may not directly cause a significant increase in driving path differences. However, such a change could serve as a valuable indicator, as altering the heading or acceleration of the ego vehicle could lead to adjustments in the driving path, potentially resulting in an increased difference between paths.
To implement this behavior feedback, the ego’s behavior is collected from its waypoints in the scenario observation $O(s)$, denoted as $\mathbf{X}_{s}=((\theta_{0},v_{0},a_{0}),(\theta_{1},v_{1},a_{1}),\ldots,(\theta_{k},v_{k},a_{k}))$, where $\theta_{i}$, $v_{i}$ and $a_{i}$ represent the heading, velocity and acceleration at timestamp $i$, respectively. To compare the behavior differences between two scenarios $s^{*}$ and $s^{\prime}$, the Maximum Mean Discrepancy (MMD) is used as the measure of distance between their behavior distributions. The MMD is a widely used statistical metric for comparing distributions and can effectively quantify differences between two sets of data. The behavior feedback function $f_{b}(s^{*},s^{\prime})$ is defined as:
$f_{b}(s^{*},s^{\prime})=f_{MMD}(\mathbf{X}_{s^{*}},\mathbf{X}_{s^{\prime}}).$ (8)
Finally, our fitness score is calculated as:
$\mathcal{D}(\tau(s^{*}),\tau(s^{\prime}))=f_{p}(s^{*},s^{\prime})+f_{b}(s^{*},s^{\prime}).$ (9)
V EMPIRICAL EVALUATION
@@
Environment. We implement Decictor on Baidu Apollo 7.0 [22] and its built-in simulation environment SimControl. Baidu Apollo 7.0 is an open-source and industrial-level ADS that supports a wide variety of driving supports.
@@
Driving Scenarios. We evaluate Decictor on a real-world Map Sunnyvale Loop, provided as part of Baidu Apollo. Following the existing works [8, 9], we use six representative scenarios and build the initial ODSs, the inputs of Algorithm 1. These scenarios have been widely tested in previous literature [8, 9, 28, 36, 10, 15, 14, 13, 17] and have covered all possible road types in the Apollo map library. The selected ODSs include left turn (S1), right turn (S2), lane following (S3), U-turn (S4), crossing intersection (S5), and existing driving road (S6). The optimality of seed ODSs is ensured through manual verification.
@@
Baselines. Since this is the first work to specifically evaluate the optimality of PPDs, and there are no baselines that are directly related to PPD testing. Thus, we design two random strategies for fair comparison with our Decictor, i.e., Random and Random-$\delta$. Specifically, 1) Random, which does not have any feedback and mutation constraints. It randomly generates scenarios by adding or removing dynamic vehicles or static obstacles; 2) Random-$\delta$, which does not have feedback but uses our non-invasive mutation $\delta$. Besides, we selected seven safety-related baselines for comparison: AVFuzzer [8], SAMOTA [10], BehAVExplor [9], DriveFuzz [13], scenoRITA [14], DoppelTest [15] and DeepCollision [17]. We understand that comparisons with the seven baselines, primarily designed for identifying safety-critical violations, may not be absolutely fair. However, our empirical results demonstrate that they can still generate scenarios with non-optimal PPDs, which would be overlooked without a consistency check. Thus, the comparison with these tools still underscore the need to test the optimality of PPDs.
Note that not all of baselines were originally evaluated in the same simulation environment as Decictor (i.e., SimControl + Apollo). Specifically, AVFuzzer, BehAVExplor, and DeepCollision were implemented based on the LGSVL simulator [37], which has been sunsetted [38]. SAMOTA and DriveFuzz are developed for Pylot [27] and Autoware [26], respectively, both running with the CARLA simulator [39]. Therefore, we invested significant effort in migrating them to our simulation environment for comparison. Since their core algorithms (e.g., surrogate models, seed selection, and feedback) are often general, we kept them the same as the original implementation. We mainly modified the simulation interface (changing the simulator APIs) so that they can run on SimControl and Apollo. Furthermore, we set the ego task consistent with our approach to facilitate comparison. Details are available on [40].
Metrics. To facilitate comparison with the baseline techniques, we have incorporated our consistency check into them to gather the generated NoDSs. In our experiments, we utilize the metrics #NoDS and #NoDS-Hum to assess the effectiveness of NoDS generation. #NoDS quantifies the number of potential NoDSs, identified through our consistency checking with a predetermined threshold $\varepsilon$.
As discussed in Section IV-A, although our mutation strategy is conservative, it is still impossible to guarantee with absolute certainty that mutations will not influence ego behavior. Additionally, the ADS may plan an alternative route that, from a human perspective, could also be deemed an optimal decision. This bears resemblance to the generation of invalid inputs in existing adversarial attacks [21] or deep learning testing methodologies [41], despite there are some mutation constraints. Considering these factors, we introduce #NoDS-Hum, representing the number of NoDSs validated by human evaluation. Specifically, the co-authors verify whether the new paths in NoDSs are truly less optimal than the paths in the seed ODSs. A NoDS is only counted in #NoDS-Hum if all authors unanimously recognize it as non-optimal, indicating a clear-cut case of a non-optimal decision.
Moreover, we also assess the success rate of non-invasive mutations, denoted as %Mutation. Specifically, the evaluation involves replaying the ego vehicle in the mutated scenario alongside the optimal path of the ODS and verifying whether the motion task can be completed on this path.
@@
Implementation. Following AVFuzzer [8], we set the population size $N$ in Decictor to $4$. In our mutation, we consider both dynamic vehicles and static obstacles (i.e., traffic cones). According to our preliminary study [42], we set the time step $\Delta t$ in the non-invasive mutation to 2 seconds and the equality threshold in the Consistency Check to $\varepsilon=0.6$. The grid size is set to 2 meters to recognize when the entire vehicle deviates from the optimal path. Similar to previous works [10, 43], we repeat each experiment 10 times to mitigate the influence of the non-determinism inherent in the simulation-based execution of the ADS. For each run, the budget is set as four hours.
V-A RQ1: Effectiveness of Decictor
TABLE I: Comparison results with baselines.
Method #NoDS (#NoDS-Hum)\,$\uparrow$ %Mutation\,$\uparrow$
S1 S2 S3 S4 S5 S6 Sum S1 S2 S3 S4 S5 S6 Avg.
AVFuzzer 0.6 (0.3) 1.5 (1.0) 0.0 (0.0) 1.2 (0.5) 1.4 (0.9) 5.0 (0.6) 9.7 (3.3) 41.5 40.5 70.6 10.6 6.7 77.5 41.2
SAMOTA 1.5 (0.9) 0.4 (0.3) 0.0 (0.0) 3.3 (1.2) 7.2 (6.2) 4.0 (0.8) 16.4 (9.4) 58.6 50.1 49.6 55.7 54.1 68.7 56.1
BehAVExplor 2.5 (1.7) 1.6 (0.7) 0.0 (0.0) 6.3 (2.3) 3.8 (3.3) 3.0 (0.4) 17.2 (8.4) 29.7 36.3 36.6 32.1 16.4 40.8 32.0
DriveFuzz 0.1 (0.0) 5.8 (2.6) 0.0 (0.0) 4.2 (1.2) 1.5 (0.9) 4.0 (0.5) 15.6 (5.2) 23.8 55.5 41.2 34.7 17.1 58.5 38.5
scenoRITA 0.7 (0.3) 0.0 (0.0) 0.0 (0.0) 0.9 (0.2) 5.2 (2.7) 2.2 (0.4) 9.0 (3.6) 92.4 94.7 98.2 94.7 96.1 96.9 95.5
DoppelTest 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 1.8 (0.6) 0.0 (0.0) 0.0 (0.0) 1.8 (0.6) 94.7 79.1 97.9 97.6 96.1 97.6 93.8
DeepCollision 2.0 (1.0) 6.1 (1.8) 0.0 (0.0) 3.2 (1.0) 2.1 (0.8) 1.9 (0.2) 15.3 (4.8) 58.0 65.5 44.0 48.4 54.9 62.0 55.5
Random 0.3 (0.0) 3.0 (1.5) 0.2 (0.2) 1.9 (0.7) 7.6 (6.8) 3.7 (0.3) 16.7 (9.5) 73.2 73.1 70.2 76.7 65.3 66.0 70.8
Random-$\delta$ 3.0 (1.4) 5.1 (4.6) 4.4 (4.4) 7.2 (2.0) 10.5 (8.1) 5.2 (0.4) 35.4 (20.9) 96.4 97.4 99.0 96.0 95.6 98.1 97.1
Decictor 9.4 (3.9) 11.9 (8.6) 9.0 (9.0) 15.7 (6.7) 12.1 (9.4) 5.8 (1.1) 63.9 (38.7) 98.0 97.5 99.4 97.5 96.6 98.2 97.9
Figure 5: Comparison of #NoDS for different tools.
V-A1 Comparative Results
Table I compares #NoDS, #NoDS-Hum, and %Mutation across six initial ODSs: S1, S2, S3, S4, S5 and S6, and Fig. 5 illustrates the distribution of #NoDS across ten iterations of each method in every scenario, where the median and the average are represented by an orange bar and a green triangle, respectively. From the results, we can find that Decictor outperforms the baselines in #NoDS and #NoDS-Hum. In detail, on average of #NoDS, Decictor outperforms the best baseline (i.e., Random-$\delta$): S1 (9.4 vs. 3.0), S2 (11.9 vs. 5.1), S3 (9.0 vs. 4.4), S4 (15.7 vs. 7.2), S5 (12.1 vs. 10.5) and S6 (5.8 vs. 5.2). Among all the detected NoDSs by Decictor, we finally identified 3.9, 8.6, 9.0, 6.7, 9.4 and 1.1 NoDS-Hum in S1, S2, S3, S4, S5 and S6 respectively. The detailed human verification results for NoDS-Hum are given on our website [40]. Even with very rigorous manual filtering, Decictor still significantly outperforms the best baseline (Random-$\delta$).
The results indicate the effectiveness of Decictor in identifying NoDSs and NoDS-Hum. We further observed that Random-$\delta$ outperforms both the Random (35.4 vs. 16.7 for the total number of NoDSs) and the best safety-oriented baseline BehAVExplor (35.4 vs. 17.2 in Sum of #NoDS). This is attributed to the effectiveness of the non-invasive mutation used in Random-$\delta$ and Decictor. It is worth noting that BehAVExplor performs better than the other safety-guided baselines due to its diversity feedback mechanism (with diverse explorations), which makes it more likely to generate NoDSs. However, scenoRITA and DoppelTest exhibit lower performance. This occurs because they search for mutations in a wider road area than others, resulting in many mutations that do not impact the ego vehicle’s motion. scenoRITA outperforms DoppelTest as it incorporates a larger number of obstacles in its mutation processes, increasing the probability of generating NoDSs. We also observed that all safety-oriented baselines fail to perform on S3. This is primarily due to their lack of consideration for the impact of static obstacles during mutation, which is the key reason for NoDSs in S3. #NoDS-Hum is lower than #NoDS due to the very strict criteria set for determining optimality.
Regarding %Mutation (Avg. column), Table I reveals that Random and safety-oriented baselines produce a maximum of 95.5% valid mutations (i.e., the original optimal path is not affected), which is lower than non-invasive mutation methods (97.1% for Random-$\delta$ and 97.9% for Decictor on average). The high rate of valid mutations demonstrates that our non-invasive mutation effectively ensures the validity of mutated objects. Note that compared with other safety-oriented baselines, scenoRITA and DoppelTest generate a higher number of valid mutations. This is because they explore a broader mutation space, most of which is disjoint from the motion space of the ego vehicle, which is less likely to conflict with the optimal path of the ego vehicle. Compared to Random-$\delta$, the average 0.8% increase of Decictor in %Mutation indicates that integrating the feedback mechanism not only maintained the effectiveness of our mutation process but also increased the detection of NoDSs.